OpenStudy (mindblast3r):

A sports drink requires two scoops of powder for every 500 mL of water. What is the consequence of mixing 10 scoops of powder with 5 L of water?

OpenStudy (mindblast3r):

like will the water be to concentrated or will it be to watery?

OpenStudy (mindblast3r):

\[2:500, 10:5000\]

OpenStudy (mindblast3r):

something like this?

OpenStudy (mindblast3r):

if I do 2/500 = 0.004. What does the 0.004 mean??

OpenStudy (campbell_st):

ok.. so the ratio is 2/500 = 10/x then x = (10 x 500)/2 which means that for 10 scoops... you need 2500 ml if you use 5000ml then it will be diluted to 1/2 strength hope it helps

OpenStudy (rational):

2 : 1/2 <=> 4 : 1 10 : 5 <=> 2 : 1

OpenStudy (rational):

*Looks the concentration is getting HALVED with "10 scoops + 5L water"

OpenStudy (mindblast3r):

who wants medal?

OpenStudy (mindblast3r):

explain dis plz "2 : 1/2 <=> 4 : 1" 10 : 5 <=> 2 : 1"

OpenStudy (mindblast3r):

how you turn 500 ml to 1/2?

OpenStudy (mindblast3r):

and 4:1??

OpenStudy (rational):

500ml = 1/2 L

OpenStudy (mindblast3r):

oh true

OpenStudy (rational):

`A sports drink requires "two" scoops of powder for every "500 mL" of water` translates to : 2 : 1/2

OpenStudy (rational):

which is same as : 4 : 1

OpenStudy (rational):

we can multiply ratios by numbers - they wont be affected

OpenStudy (mindblast3r):

but written as a fraction it would be 2/500?

OpenStudy (rational):

yes : \(\dfrac{2}{500ml} = \dfrac{2}{\frac{1}{2}L}\)

OpenStudy (rational):

let me rephrase the problem and give u

OpenStudy (mindblast3r):

when i put it on the calculator it gets different answers doh

OpenStudy (rational):

A sports drink requires two scoops of powder for every \(\color{red}{\dfrac{1}{2}L}\) of water. What is the consequence of mixing 10 scoops of powder with 5 L of water?

OpenStudy (campbell_st):

just keep things in mL its the given measurements... don't confuse things by changing the given measurements.

OpenStudy (rational):

you need to change everything into mL or L, its your choice ! btw you're given things in mL alone

OpenStudy (campbell_st):

just know 5000 ml = 5L however either method gives the same answer... as long as the units of capacity are the same

OpenStudy (rational):

*btw you're NOT given things in mL alone

OpenStudy (mindblast3r):

this is what i'd do, 2/500 = 0.004 10/5000 = 0.002

OpenStudy (mindblast3r):

but what the heck does that answer even mean??

OpenStudy (mindblast3r):

am i even doing it right?

OpenStudy (rational):

you're doing it right, since u are using mL, the answer resulted in decimals and it is becoming hard to interpret

OpenStudy (mindblast3r):

ahh i get it.

OpenStudy (rational):

2/500 = 0.004 10/5000 = 0.002 just means this : the ratio has decreased by a factor of 2

OpenStudy (mindblast3r):

what would be the easiest method then?

OpenStudy (rational):

above should make sense after scratching ur head 2-3 times ?

OpenStudy (mindblast3r):

": the ratio has decreased by a factor of 2" wut mean?

OpenStudy (rational):

0.002 = 0.004/2 so clearly the ratio has \(\cdots \) fill in whichever makes sense to u

OpenStudy (mindblast3r):

i mean what unit of measurement is the 0.002?

OpenStudy (rational):

10/5000 = 0.002 10 scoops per 5000mL of water so the unit is scoops/mL

OpenStudy (rational):

which is clearly a nonsense unit

OpenStudy (mindblast3r):

rofl

OpenStudy (rational):

change it to L and work it agian, you will ask less questions then :P

OpenStudy (mindblast3r):

could you do it for me in L please?

OpenStudy (rational):

A sports drink requires two scoops of powder for every 500 mL of water. What is the consequence of mixing 10 scoops of powder with 5 L of water?

OpenStudy (rational):

\(two ~scoops\) for every \(500ml \) that means how many scoops for every \(L\) ?

OpenStudy (mindblast3r):

1000 ml = 1L 500 = 1/2 L

OpenStudy (rational):

yes \(two ~scoops\) are there for every \(1/2L\) of water so that makes \(four ~scoops \) for every \(1L\) of water

OpenStudy (rational):

the ratio in \(L\) is : \(4 : 1\)

OpenStudy (rational):

if that makes any sense...

OpenStudy (mindblast3r):

ah you just increased the ratio proportionally.

OpenStudy (mindblast3r):

by 2

OpenStudy (mindblast3r):

doubled it

OpenStudy (rational):

exactly !

OpenStudy (rational):

so the required ratio is \(4 : 1\) (four scoops for every 1L of water)

OpenStudy (rational):

does that look pleasing... to interpret ?

OpenStudy (mindblast3r):

yes.

OpenStudy (mindblast3r):

ahhh so so increasing the ratio proportionally makes it easier to interpret?

OpenStudy (rational):

yes somehow our brain catches natural numbers like 1,2,3,4... more quickly than bfractions/decimals lol

OpenStudy (mindblast3r):

lololol your dispaly pic is so funny.

OpenStudy (mindblast3r):

anyway lets go back to the 0.002, what does this mean exactly?

OpenStudy (rational):

0.002 scoops per mL of water

OpenStudy (mindblast3r):

1 = 1 school?

OpenStudy (mindblast3r):

scoop*

OpenStudy (mindblast3r):

rofl

OpenStudy (mindblast3r):

like if you add .998 to 0.002 = 1 scoop?

OpenStudy (rational):

its like measuring ur height with a kilometer scale

OpenStudy (rational):

my height is 0.0016KM interpret that^

OpenStudy (mindblast3r):

roflcopter lolol i get it.

OpenStudy (rational):

which is technically okay, but it hurts ur head very badly u will never knw that easily whats my height is

OpenStudy (mindblast3r):

i mean i get what your trying to say, not your height in kilometers lol

OpenStudy (mindblast3r):

ya... with that number i can't imagine how tall you are. But if you told me something like your 6ft i'd definitely know how tall you are.

OpenStudy (rational):

exactly ! so lets change `0.002 scoops per mL` to something practical units

OpenStudy (rational):

0.002 scoops per mL 0.002*1000 scopps per 1L 2 scoops per 1L

OpenStudy (rational):

now it should be easy to interpret ^

OpenStudy (rational):

0.002 scoops per mL is same as saying : 2 scoops per 1Liter

OpenStudy (mindblast3r):

wtf? that's wrong 2 scoops per 500 ml 2:1/2 = 4:1

OpenStudy (mindblast3r):

it's 4 scoops per liter lol

OpenStudy (rational):

4 scoops per liter is the requirement

OpenStudy (rational):

2 scoops per liter is what u get by combining `5 scoops with 10 liters of water`

OpenStudy (rational):

2/500 = 0.004 10/5000 = 0.002

OpenStudy (mindblast3r):

"2 scoops per liter is what u get by combining 5 scoops with 10 liters of water" Are you asking me or telling me?

OpenStudy (rational):

telling you :/

OpenStudy (mindblast3r):

wtf?

OpenStudy (mindblast3r):

wouldn't it be the same answer then?

OpenStudy (mindblast3r):

2/500 = 0.004 10/5000 = 0.002

OpenStudy (mindblast3r):

lol i'm really sorry, i'm trying to get this.

OpenStudy (mindblast3r):

do you think i'm over thinking this?

OpenStudy (rational):

nope, u need to start over and work it purely in liters

OpenStudy (rational):

and u need to forget about 0.002 and mL completely

OpenStudy (mindblast3r):

okay then so I'll write the out the answer to the question now

OpenStudy (mindblast3r):

and you tell me wut you think

OpenStudy (rational):

my first reply in this thread has everything thats needed, but im not sure u understood it...

OpenStudy (mindblast3r):

\[\frac{ 2 }{ 0.5 } \times \frac{ 2 }{ 2} = \frac{ 4 }{ 1 }\]

OpenStudy (mindblast3r):

i just doubled the recipe lol

OpenStudy (rational):

looks good^

OpenStudy (mindblast3r):

\[\frac{ 10 }{ 5 } = 2 scoops/L\]

OpenStudy (mindblast3r):

\[\frac{ 4 }{ 1}= 4scoops/L\]

OpenStudy (rational):

yes, compare them both

OpenStudy (rational):

which one has more scoops per L ?

OpenStudy (mindblast3r):

the 4/1 one